# RHINO: A large horn antenna for detecting the 21cm global signalRHINO: A large horn antenna for detecting the 21cm global signal

### v1

@@ 4.2 Warm calibration loads
where $\Delta T=T_{\rm rec}^{\rm phys}-T_{\rm calsrc}^{\rm phys}$ is the difference between the physical temperatures of the lossy cable at the receiver input port and the calibration source, and $L$ is the insertion loss of the cable in dB. By using short lengths of semi-rigid coaxial cable with SMA connectors, we hope to minimise the insertion loss to below $L\lesssim 0.3$ dB. For a calibration load at 383 K, receiver port at 303 K, and insertion loss of 0.3 dB, the difference between the physical and effective temperature of the calibration source would then be $\approx-2.7$ K, or about a 0.7% correction to the calibration source power level. Stabilising the temperature of the cables/receiver, as well as the calibration source, is therefore necessary to avoid fluctuations in the reference power level at the sub-percent level. Note that the effect of the cables also includes reflections due to inexact matching, as well as signal attenuation and added noise power (Zannoni et al., 2008).
@@ 5 Performance forecasts
Figure 10: Posterior predictive distributions for the global signal as predicted by 21cmVAE, for the RHINO ($60-85$ MHz) and wideband ($60-150$ MHz) scenarios, for simulated data with perfect calibration, no RFI flags, and thermal noise equivalent to 20 full nights of observing. To permit a fair comparison across the full frequency range, a frequency-dependent Airy beam for a 7m aperture was assumed. Fits are shown to simulated data without foregrounds, and with PCA-subtracted foregrounds (12 modes removed). The median is shown as a solid coloured line, and the 68% and 95% credible regions are shown as shaded bands. The true (input) global signal is shown as a solid grey line with time-averaged noise included.
@@ 5.1 Basic simulations and model fitting
Our focus here is on recovery of physical model parameters for ‘standard’ models that do not attempt to explain the EDGES signal. We use the 21cmVAE emulator (Bye et al., 2022) to define a broad space of possible global signal models, parametrised by 7 values that summarise pertinent physical processes, such as the star formation efficiency, $f_{*}$, X-ray heating efficiency, $f_{\rm X}$, the minimum circular velocity for a halo to support star formation, $V_{c}$, and the low-energy cut-off of the X-ray SED, $\nu_{\rm min}$ (Cohen et al., 2020). Most of the variation in the global signal can be captured by these four parameters; the other 3 parameters (ionising photon mean free path, $R_{\rm mfp}$, optical depth to the CMB, $\tau$, and X-ray SED power-law slope, $\alpha$) have a relatively minor effect on the global signal (Bye et al., 2022).
@@
Fig. 9 shows 80 randomly-chosen examples of plausible models produced by the emulator. The parameter values were randomly drawn from uniform distributions with ranges at or close to the maximum ranges of the 21cmVAE emulator, i.e. $f_{*}\in[10^{-4},0.5]$, $V_{c}\in[4.2,100]$ km/s, $f_{X}\in[10^{-5},100]$, and $\nu_{\rm min}\in[0.1,3.0]$ keV, with the other parameters fixed to values of $\tau=0.07$, $\alpha=1.25$, and $R_{\rm mfp}=29.6$ Mpc, as in Bye et al. (2022). The RHINO band is highlighted in red, and while it can be seen that some models have narrower absorption features that fall within this region, many others have broader features, leaving only a relatively smooth and featureless gradient across the RHINO band. The smoother the global signal is, the harder it is likely to be to disentangle it from similarly smooth foregrounds and systematic artefacts. Note that the emulator predictions for the global signal are noisy at small frequency separations, and so we apply a Savitzky-Golay filter (Savitzky & Golay, 1964) to smooth out the small-scale variation, which would otherwise be confused with thermal noise.
@@
Next, we chose one of the randomly generated models that has an absorption feature roughly 100 mK deep within the RHINO band (see Fig. 10). This was added to basic simulations that used a frequency-dependent Airy beam pattern for a 7.2m aperture telescope, a sky model containing only diffuse emission from GSM 2016 (Zheng et al., 2017), and thermal noise consistent with 20 nights of observing with short drift scans of 72 minutes with 10s integrations. Instrumental systematics and calibration errors were not included; while these are important effects, we only seek to understand the impact of the RHINO bandwidth on simple idealised scenarios in this study.
@@
With the simulations in hand, we performed a joint foreground plus 21cm global signal model fit. The 21cm models were provided by 21cmVAE, with 4 free parameters as discussed above. The foregrounds were modelled using an idealised Principal Component Analysis (PCA) technique, which we implemented by taking the simulated foreground model (including beam effects, but without noise) for the relevant frequency window for each scenario and calculating the first few principal components of its frequency-frequency covariance matrix (constructed by averaging over time). The first 12 PCA modes were used to define the foreground model that is included in the fits. These modes had significant smooth spectral variation based on visual inspection, whereas subsequent modes became more noise-like. The PCA mode amplitudes and 4 global signal model parameters were then jointly sampled using emcee (Foreman-Mackey et al., 2013), with uniform priors on all parameters. For simplicity, we worked with the time-averaged data for the model fitting, although one could in principle attempt to fit the foreground separately for each time sample.
@@ 5.2 Global signal recovery
Both of these cases demonstrate the need for robust physical modelling of the instrument and sky signals. Blind (data-driven) foreground removal methods permit too much flexibility in the foreground model, resulting in most of the 21cm global signal being absorbed by the foreground modes. This results in strong correlations between the foreground and global signal parameters, making it difficult to unambiguously disentangle them. Physical modelling (e.g. Sathyanarayana Rao et al., 2017) should result in a much more constrained range of possible foreground behaviours, at the expense of significantly increased model complexity. From the results shown in Fig. 10, we can conclude that much more constrained foreground models will be necessary if RHINO is to be able to recover the global signal – the limited frequency range increases the confusion between foregrounds and signal. Similar measures would be needed even for an experiment with much wider bandwidth, however, as strong correlations between foreground modes and the smooth global signal persist, resulting in a biased inference in the particular example shown above.
@@ 6 Prototype system
Figure 14: Sky temperature as a function of LST and frequency, estimated from the GSM 2016 sky model smoothed by a Gaussian beam with FWHM $\approx\lambda/D$ for $D=7$m. The assumed site is Jodrell Bank Observatory.
@@
The prototype antenna is a scaled-down pyramidal horn with a wooden frame and welded mesh conductive surface consisting of multiple overlapping segments. It is attached to a rectangular waveguide of similar construction, and targets a centre frequency of $\sim 350$ MHz. The right panel of Fig. 12 shows the flare section during construction at JBO. The frame is made in several trapezoidal sections with reinforcing ribs inside, and joined together with metal plates and brackets. Inside the frame are multiple sections of 13mm galvanised welded mesh, stapled to the wooden frame, with several cm overlap between each section. The structure is positioned to point vertically at the zenith, with metal scaffolding poles used to support it. At the aperture, the conducting surface has a square profile of side length 2.19m, giving $\lambda/D\approx 22^{\circ}$ at 350 MHz. The purpose of the scaled-down prototype is to experiment with fabrication and beam measurement methods, to study the impact of environmental conditions (e.g. thermal expansion, wind loading, and soil moisture) on the real-world antenna pattern (c.f. Pattison et al., 2024a).

### v2

@@ 4.2 Warm calibration loads
where $\Delta T=T_{\rm rec}^{\rm phys}-T_{\rm calsrc}^{\rm phys}$ is the difference between the physical temperatures of the lossy cable at the receiver input port and the calibration source, and $L$ is the insertion loss of the cable in dB. By using short lengths of semi-rigid coaxial cable with SMA connectors, we hope to minimise the insertion loss to below $L\lesssim 0.3$ dB. For a calibration load at 383 K, receiver port at 303 K, and insertion loss of 0.3 dB, the difference between the physical and effective temperature of the calibration source would then be $\approx-2.7$ K, or about a 0.7% correction to the calibration source power level. Stabilising the temperature of the cables/receiver, as well as the calibration source, is therefore necessary to avoid fluctuations in the reference power level at the sub-percent level. Note that the effect of the cables also includes reflections due to inexact matching, as well as signal attenuation and added noise power (2008ApJ...688...12Z).
The choice of calibration load temperatures is discussed in Sect. 4.3. A substantially higher temperature is preferable for the hottest load. This presents a number of challenges in terms of temperature control, safety, and the materials used however; the PTFE dielectric commonly used in coaxial cable has a melting point of around 600 K for example. Instead, it may be more practical to use a non-thermal reference source instead, such as a noise diode. These can achieve much larger effective temperatures, with good temporal stability, but would need an additional bandpass calibration of the source to be performed.
4.3 Optimisation of observing time/calibration cycles
In this section, we discuss an optimised way of choosing how much time should be spent observing each calibration load vs the antenna.
Pairs of well-characterised reference sources can be used to calibrate a given source by forming a quotient
$q=\frac{P_{\text{src}}-P_{\text{L}}}{P_{\text{NS}}-P_{\text{L}}},$ (19)
where $P_{\text{src}}$ is a measurement of the power spectral density (PSD) for a source input to the receiver, and $P_{\text{NS}}$ and $P_{\text{L}}$ are PSDs from measurements of two different reference noise-sources of known noise temperature. Such quotients can then be used to fit gains, noise-wave parameters, and other degrees of freedom needed to calibrate the noise temperatures of an input source such as an antenna (2012RaSc...47.0K06R; 2021MNRAS.505.2638R; 2022MNRAS.517.2264M).
Eq. 19 is a first-order approximation that assumes that the two reference noise sources are perfectly impedance-matched to the receiver. This removes dependence on signal gains that occur after the switching interface, as well as the self-generated noise power from the receiver that does not undergo reflections before being measured. Additionally, the assumption is made that the PSD measurements are taken within a time frame in which the receiver gains are constant. This is equivalent to keeping the total integration time under the knee-frequency of the receiver so that the PSD measurements are dominated by Gaussian thermal noise.
The signal-to-noise ratio (SNR) when measuring $q$ limits the precision to which the receiver noise-wave parameters can be measured, which in turn affects the calibrated value of the sky temperature, $T_{\text{sky}}$. It is therefore beneficial to maximise this SNR if possible. This can be achieved by tuning the integration times for the PSD measurements as well as the choice of calibrator noise temperatures.
The optimal observing strategy is well-known in Dicke-switched systems, for example (dicke1946measurement). Observing time is split equally between the source and calibrator, and when $T_{\rm src}\approx T_{\rm L}$, it can be shown that the effects of gain fluctuations are cancelled out when observing a source and calibrator of the same radiometric power.
We now present a derivation of an optimised calibration strategy for our receiver system, subject to the assumptions above. Assuming that the three PSD measurements for a single estimate of $q$ are completed in a total time $\tau$, with a channel bandwidth $\Delta\nu$, the radiometer equation can be used to derive the variance for each measurement based on the total system noise temperature. The variance of a single frequency channel for a PSD measurement of a given source, $i$, can be expressed as
$\sigma_{P_{i}}^{2}=\left(gk_{\rm B}T_{\text{sys}}\right)^{2}\frac{\Delta\nu}{\phi_{i}\tau}=\frac{P_{i}^{2}}{\Delta\nu\phi_{i}\tau},$ (20)
where $g$ is the receiver gain, $k_{\rm B}$ is Boltzmann’s constant, $T_{\text{sys}}$ is the receiver system temperature and $\phi_{i}$ is the fraction of the total time $\tau$ spent on the single PSD measurement such that $\phi_{\text{src}}+\phi_{\text{L}}+\phi_{\text{NS}}=1$.
Neglecting gain fluctuations and spectral leakage between channels in this analysis (which introduce correlations in frequency and time), when the error on the PSDs are propagated in quadrature, the fractional variance for a measurement of $q$ can be expressed as
$\left(\frac{\sigma_{q}}{q}\right)^{2}=\frac{1}{\Delta\nu\tau}\left[\frac{P_{\rm src}^{2}}{(P_{src}-P_{\rm L})^{2}}\frac{1}{\phi_{\rm src}}\right.+\frac{P_{\rm NS}^{2}}{(P_{\rm NS}-P_{\rm L})^{2}}\frac{1}{\phi_{\rm NS}}$
$+\left.\frac{P_{\rm L}^{2}\left(P_{\rm src}-P_{\rm NS}\right)^{2}}{\left(P_{\rm NS}-P_{\rm L}\right)^{2}\left(P_{\rm src}-P_{\rm L}\right)^{2}}\frac{1}{\phi_{\rm L}}\right],$ (21)
where $P_{\rm src}$, $P_{\rm NS}$ and $P_{\rm L}$ retain the definitions defined in Eq. 19. This can be expressed in terms of noise temperatures and reflection coefficients for the receiver and source by substituting for Eq. 6.
Figure 10: The predicted fractional standard deviation for a single measurement of $q$ when $T_{\rm L}=$ 380 K, $T_{\rm NS}=$ 490 K, the total observing time $\tau=$45 s and $\delta\nu=$1.22 kHz. The bottom plot shows the ratio between the standard deviations of $q$ for the optimised equal strategy and the optimised observation strategy.
Eq. 4.3 can be minimised with respect to the fractional observation times $\phi_{i}$ subject to the constraint $\phi_{\text{src}}+\phi_{\text{L}}+\phi_{\text{NS}}=1$, to yield
$\hat{\phi_{i}}=\alpha m_{i}c_{i}$ (22)
where
$m_{\rm src}$ $=$ $\frac{P_{\rm src}}{P_{\rm src}-P_{\rm L}},$ (23)
$m_{\rm NS}$ $=$ $\frac{P_{\rm NS}}{P_{\rm NS}-P_{\rm L}},$ (24)
$m_{\rm L}$ $=$ $\frac{P_{\rm L}\left(P_{\rm src}-P_{\rm NS}\right)}{\left(P_{\rm NS}-P_{\rm L}\right)\left(P_{\rm src}-P_{\rm L}\right)},$ (25)
$\alpha$ $=$ $\left(m_{\rm src}+c_{\rm NS}m_{\rm NS}+c_{\rm L}m_{\rm L}\right)^{-1}$ (26)
and
$c_{\rm src},c_{\rm NS},c_{\rm L}=\begin{cases}+1,-1,-1&\text{if }P_{\rm src}\leq P_{\rm L}\\ +1,+1,-1&\text{if }P_{\rm L}<P_{\rm src}\leq P_{\rm NS}\\ +1,+1,+1&\text{if }P_{\rm NS}<P_{\rm src}.\end{cases}$
Fig. 10 shows a comparison between the optimal fractional observation times and the case in which equal time is spent on each target ($\phi_{\text{src}}=\phi_{\text{L}}=\phi_{\text{NS}}$) as a function of $T_{\rm src}$. Here, $P_{\rm L}$ and $P_{\rm NS}$ take values sourced from the heated impedance-matched loads outlined in Sect. 3.2 ($T_{\rm L}=$ 380 K and $T_{\rm NS}=$ 490 K). The channel bandwidth, $\delta\nu$, is set to be 1.22 kHz, the total observing time is set to $\tau=$ 45 s and the gain is set to unity. $P_{\rm src}$ is calculated assuming that $P_{\rm src}=gk_{B}\Delta\nu(T_{\rm src}+T_{0})$ where $T_{0}$ is the receiver-induced, non-reflected noise temperature (reflection terms are absorbed into $T_{\rm src}$).
In both cases there are singularities when $T_{\rm src}=T_{\rm L}$, resulting in a sharp increase in the variance of $q$ and a resulting decrease in the SNR. Although the SNR can be increased with longer integration times, in actual observations this is limited to the time scale on which the gain fluctuations dominate over the thermal white noise ($\tau\lesssim 1/f_{\rm knee}$).
Optimising the fractional observation times when $T_{\rm src}\approx T_{\rm L}$ can lead to a significant improvement in the SNR on the quotient, of a little over 20% compared with the equal-time case. When $T_{\rm src}$ becomes much larger than $T_{\rm NS}$, the improvement over the equal-time case is only around 787-8% however. This shows a relative insensitivity to the fraction of time spent observing each calibrator, implying that longer can be spent observing through the antenna without significantly degrading the calibration measurements.
As $T_{\rm src}$ may also encapsulate reflected noise power contributions, these must also be considered when attempting to improve the SNR for measurements of $q$. Therefore, an approximation of the expected noise temperature of $T_{\rm src}$, including the reflected noise wave terms using Eqs. 7 and 8, can be used to ensure a suitable difference in noise temperature from impedance-matched calibrator $T_{\rm L}$. As this depends on measurements of the reflection coefficients and receiver noise wave parameters, this tuning of $T_{\rm L}$ may follow from laboratory measurements of receiver and source properties.
To extract the noise wave parameters of the receiver, measurements of $q$ can be made for calibration standards such as long open and shorted cables, as well as other impedance mismatched calibrators, to induce known reflections and phase delays, producing spectral features that can be measured and used to derive a calibration (2021MNRAS.505.2638R; 2023arXiv230511479P). These measurements are typically made for standards at an ambient temperature $T_{\rm amb}$, with a resulting noise temperature $T_{\rm src}\approx T_{\rm amb}$. Tuning the noise temperature of $T_{\rm L}$ to be sufficiently higher than $T_{\rm amb}$ (so that the total $T_{\rm src}$, including the reflected noise wave terms, does not exceed or approach $T_{\rm L}$ for multiple calibration standards) will result in a decrease in noise on the measurements of $q$. This in turn will result in an increase in accuracy of measurements of the induced spectral features and subsequent extraction of the noise wave parameters through fits and other methods. The calibrated $T_{\rm sky}$ will then be more precise, with less contamination from the reflected receiver noise power.
In the fully optimised case, in which the highest SNR is extracted with maximal time spent observing the sky, $T_{\rm NS}$ should exceed the expected antenna noise temperature. Given the typical sky temperatures of $\gtrsim 10^{3}$ K at these frequencies, this would need to be achieved with a non-thermal source, e.g. a noise diode.
4.4 Calibration requirements
In general, the precision with which we can determine the noise wave calibration parameters will be a function of temporal and frequency scale. There is a relatively complex interaction between calibration and the foreground removal and 21cm model fitting procedures, as calibration errors can modulate these signals, producing spurious features that can bias or be confused with 21cm absorption features for example. Additive systematics can also partially mimic or obscure the cosmological signal. The way in which these interactions play out depends on the methods used to propagate the calibration, remove the foregrounds, and extract the 21cm signal – for instance, data-driven methods reliant on spectral smoothness may absorb the effects of calibration errors differently from more constrained methods based on forward modelling. As such, it is difficult to put strict calibration requirements on the system a priori, as they will depend on a variety of analysis choices that we will need to make further down the line, i.e. once the system is operational and we have a good idea of the behaviour of the receiver. Other groups have performed simulated analyses (e.g. 2022NatAs...6..984D), subject to this caveat.
Nevertheless, we have tried to give some sense of the importance of errors on different calibration parameters, and how these could obscure the 21cm signal. Fig. 11 shows the 21cm global signal for a set of example physical models, plotted with various systematic and calibration error terms. The amplitude of these terms is chosen to indicate the approximate measurement precision required to restrict each term to a typical value of around 100 mK, i.e. comparable to the expected global signal. Depending on the real-world spectral and temporal characteristics of each of these terms, they may be more or less difficult to separate from the 21cm signal, so these numbers should be taken only as a guide, but we can at least get a sense of the size of calibration errors that could be tolerated without severely affecting the 21cm signal extraction. Note that we have omitted several important potential sources of calibration error, including the CW source and the effects of soil conductivity/groundspill, as these will be the subject of much more detailed treatments in forthcoming publications.
Figure 11: Temperature difference vs frequency caused by mis-estimates of various noise wave parameters. The values have been chosen to give a temperature difference of similar size to the physical 21cm models described in Sect. 5.1 and Fig. 12, shown as faint grey lines. An EDGES-like absorption feature is also shown as a solid black line.
To derive these illustrative examples we use Eqs. 7 and 8, with EDGES-like values for the noise-wave parameters (see Extended Data Figure 5 of 2018Natur.555...67B), and the antenna reflection coefficient derived from the calculated $S_{11}$ shown in Fig. 5. All of these parameters will of course differ when the system is deployed, but these choices should be sufficiently realistic for this illustration. We then introduced a small error on each parameter in turn, setting all other parameters to their fiducial values, and then subtracted off the fiducial model to leave a temperature difference $\Delta T$ due to each term. The size of the error was chosen by hand to make the typical value of $\Delta T\sim 100$ mK, and is shown for each term in the legend of Fig. 11. For further comparison, we also plotted a simple illustrative error due to an imperfect foreground model, represented by a systematic fractional error of $0.01\%$ on the temperature of each pixel of the sky model at each frequency.
It can be seen from Fig. 11 that the calibration requirements are quite stringent, and that the calibration errors (in this simple approach) are spectrally smooth, and so do risk introducing spectral features that could be at least partially confused with the 21cm global signal. The noise wave temperature coefficients would need to be constrained to 0.1 K or better, while the antenna and receiver reflection coefficients must be measured to better than 0.01% (errors below 40-40 dB). The latter should ultimately be achievable given that the dynamic range of modern VNAs exceeds this level by tens of dB. In comparison, the simple foreground model error shown here highlights the difficulty of building sufficiently precise forward models of the sky; a mere 0.01% error is enough to produce a temperature difference of similar size to the expected 21cm signal. Notably, the ‘typical’ expected 21cm signal is smooth and slowly-varying across our band, increasing the difficulty of distinguishing it from calibration errors compared with the more distinctive EDGES signal.
@@ 5 Performance forecasts
Figure 13: Posterior predictive distributions for the global signal as predicted by 21cmVAE, for the RHINO (608560-85 MHz) and wideband (6015060-150 MHz) scenarios, for simulated data with perfect calibration, no RFI flags, and thermal noise equivalent to 20 full nights of observing. To permit a fair comparison across the full frequency range, a frequency-dependent Airy beam for a 7m aperture was assumed. Fits are shown to simulated data without foregrounds, and with PCA-subtracted foregrounds (12 modes removed). The median is shown as a solid coloured line, and the 68% and 95% credible regions are shown as shaded bands. The true (input) global signal is shown as a solid grey line with time-averaged noise included.
@@ 5.1 Basic simulations and model fitting
Our focus here is on recovery of physical model parameters for ‘standard’ models that do not attempt to explain the EDGES signal. We use the 21cmVAE emulator (2022ApJ...930...79B) to define a broad space of possible global signal models, parametrised by 7 values that summarise pertinent physical processes, such as the star formation efficiency, $f_{*}$, X-ray heating efficiency, $f_{\rm X}$, the minimum circular velocity for a halo to support star formation, $V_{c}$, and the low-energy cut-off of the X-ray SED, $\nu_{\rm min}$ (2020MNRAS.495.4845C). Most of the variation in the global signal can be captured by these four parameters; the other 3 parameters (ionising photon mean free path, $R_{\rm mfp}$, optical depth to the CMB, $\tau$, and X-ray SED power-law slope, $\alpha$) have a relatively minor effect on the global signal (2022ApJ...930...79B).
@@
Fig. 12 shows 80 randomly-chosen examples of plausible models produced by the emulator. The parameter values were randomly drawn from uniform distributions with ranges at or close to the maximum ranges of the 21cmVAE emulator, i.e. $f_{*}\in[10^{-4},0.5]$, $V_{c}\in[4.2,100]$ km/s, $f_{X}\in[10^{-5},100]$, and $\nu_{\rm min}\in[0.1,3.0]$ keV, with the other parameters fixed to values of $\tau=0.07$, $\alpha=1.25$, and $R_{\rm mfp}=29.6$ Mpc, as in 2022ApJ...930...79B. The RHINO band is highlighted in red, and while it can be seen that some models have narrower absorption features that fall within this region, many others have broader features, leaving only a relatively smooth and featureless gradient across the RHINO band. The smoother the global signal is, the harder it is likely to be to disentangle it from similarly smooth foregrounds and systematic artefacts. Note that the emulator predictions for the global signal are noisy at small frequency separations, and so we apply a Savitzky-Golay filter (1964AnaCh..36.1627S) to smooth out the small-scale variation, which would otherwise be confused with thermal noise.
@@
Next, we chose one of the randomly generated models that has an absorption feature roughly 100 mK deep within the RHINO band (see Fig. 13). This was added to basic simulations that used a frequency-dependent Airy beam pattern for a 7.2m aperture telescope, a sky model containing only diffuse emission from GSM 2016 (2017MNRAS.464.3486Z), and thermal noise consistent with 20 nights of observing with short drift scans of 72 minutes with 10s integrations. Instrumental systematics and calibration errors were not included; while these are important effects, we only seek to understand the impact of the RHINO bandwidth on simple idealised scenarios in this study.
@@
With the simulations in hand, we performed a joint foreground plus 21cm global signal model fit. The 21cm models were provided by 21cmVAE, with 4 free parameters as discussed above. The foregrounds were modelled using an idealised Principal Component Analysis (PCA) technique, which we implemented by taking the simulated foreground model (including beam effects, but without noise) for the relevant frequency window for each scenario and calculating the first few principal components of its frequency-frequency covariance matrix (constructed by averaging over time). The first 12 PCA modes were used to define the foreground model that is included in the fits. These modes had significant smooth spectral variation based on visual inspection, whereas subsequent modes became more noise-like. The PCA mode amplitudes and 4 global signal model parameters were then jointly sampled using emcee (2013PASP..125..306F), with uniform priors on all parameters. For simplicity, we worked with the time-averaged data for the model fitting, although one could in principle attempt to fit the foreground separately for each time sample.
@@ 5.2 Global signal recovery
Both of these cases demonstrate the need for robust physical modelling of the instrument and sky signals. Blind (data-driven) foreground removal methods permit too much flexibility in the foreground model, resulting in most of the 21cm global signal being absorbed by the foreground modes. This results in strong correlations between the foreground and global signal parameters, making it difficult to unambiguously disentangle them. Physical modelling (e.g. 2017AJ....153...26S) should result in a much more constrained range of possible foreground behaviours, at the expense of significantly increased model complexity. From the results shown in Fig. 13, we can conclude that much more constrained foreground models will be necessary if RHINO is to be able to recover the global signal – the limited frequency range increases the confusion between foregrounds and signal. Similar measures would be needed even for an experiment with much wider bandwidth, however, as strong correlations between foreground modes and the smooth global signal persist, resulting in a biased inference in the particular example shown above. Recent work on Bayesian sky models for 21cm global signal applications offers a promising approach to address this issue (e.g. 2025arXiv250101417C; 2025arXiv250621258I; 2025arXiv250911894N).
@@ 6 Prototype system
The prototype antenna is a scaled-down pyramidal horn with a wooden frame and welded mesh conductive surface consisting of multiple overlapping segments. It is attached to a rectangular waveguide of similar construction, and targets a centre frequency of $\sim 350$ MHz. The right panel of Fig. 15 shows the flare section during construction at JBO. The frame is made in several trapezoidal sections with reinforcing ribs inside, and joined together with metal plates and brackets. Inside the frame are multiple sections of 13mm galvanised welded mesh, stapled to the wooden frame, with several cm overlap between each section. The structure is positioned to point vertically at the zenith, with metal scaffolding poles used to support it. At the aperture, the conducting surface has a square profile of side length 2.19m, giving $\lambda/D\approx 22^{\circ}$ at 350 MHz. The purpose of the scaled-down prototype is to experiment with fabrication and beam measurement methods, to study the impact of environmental conditions (e.g. thermal expansion, wind loading, and soil moisture) on the real-world antenna pattern (c.f. 2025MNRAS.538.1301P).